# Market-Driven Equilibria for Distributed Solar Panel Investment

### v1

@@ III-B Real-time Pricing Mechanisms
Within the product-differentiated real-time market, the traded product is solar electricity. Each seller $i\in\mathcal{I}_{\mathrm{s}}$ can sell any quantity of solar generation within the set $\mathcal{Q}_{i}^{\mathrm{s}}=[0,cG]$, which depends on the realized value of $G$. The seller payoff function is
$\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)=\pi q^{\mathrm{s}}_{i},$
where market equilibrium price $\pi$ and cleared quantity $q^{\mathrm{s}}_{i}$ will be defined momentarily. Given the inelastic load $L$, a buyer $i\in\mathcal{I}_{\mathrm{b}}$ can purchase any quantity of solar within the range $\mathcal{Q}^{\mathrm{b}}_{i}=[0,L]$. Any remaining demand not met by solar will be fulfilled by the utility backstop at the price $\pi_{\mathrm{u}}$. The buyer payoff function is
$\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=(v_{i}-\pi)q^{\mathrm{b}}_{i}-\pi_{\mathrm{u}}(L-q^{\mathrm{b}}_{i}),$
where $q_{i}^{\mathrm{b}}$ is the cleared quantity for buyer $i$. Denote $q^{\mathrm{s}}:=\{q^{\mathrm{s}}_{i}\}_{i\in\mathcal{I}_{\mathrm{s}}}$ and $q^{\mathrm{b}}:=\{q^{\mathrm{b}}_{i}\}_{i\in\mathcal{I}_{\mathrm{b}}}$.
We use the following equilibrium notion to characterize the market outcome:
@@
We intentionally omit the price discovery process and focus on the CE, as our interest lies in the outcomes rather than the mechanism used to obtain them. Such outcomes can arise, for instance, from buyers and sellers submitting bids to a market operator who clears the market. In our non-atomic game model, where all agents are infinitesimally small and act as price takers, it can be shown that an NE of such a bidding process corresponds to a CE that we consider here. Detailed models of price formation have been extensively studied in prior work (see references in Section I) and are beyond the scope of this paper.
With the equilibrium price $\pi$ and the cleared quantities $q^{\mathrm{s}}_{i}$ for sellers in the product-differentiated real-time market, the total expected revenue for sellers with an aggregate panel capacity $c$ over the panels’ lifespan can be calculated as
@@ III-C Contract-based Market
where $(z)_{+}:=\max(z,0)$, and $q_{i}^{\mathrm{s}}$ and $q_{i}^{\mathrm{b}}$ are the cleared quantities (i.e., panel capacities) for the seller and buyer, respectively.
@@
The proposed market model serves as a general framework capable of representing a variety of solar market mechanisms, provided that the traded product and the payoff structures for buyers and sellers are appropriately defined. For example, both real-time pricing and contract-based arrangements can be formulated within this framework, with key distinctions captured through the specification of the traded product, the feasible sets of buyers and sellers, and the corresponding payoff functions. We abstract away additional implementation-specific features, as our primary objective is to investigate how different market structures shape long-term investment outcomes. While these mechanisms may differ in operational frequency, often leading to variations in price volatility as discussed in prior work [11, 12, 13], such factors do not directly affect the investment equilibrium when only expected returns matter.
@@
IV Competitive Equilibria of Solar Markets
In this section, we derive the expected revenue for panel owners as a function of panel capacity $c$.
@@ IV Competitive Equilibria of Solar Markets
A CE for product-differentiated real-time market (prt) and single-product real-time market (srt), are listed in Table I, where $\overline{F}_{V}(\cdot):=1-F_{V}(\cdot)$ is the complementary cumulative distribution function of $v_{i}$, and $\overline{\pi}^{\mathrm{b}}_{i}:=\pi_{\mathrm{u}}+v_{i}$. Furthermore, the expected total revenue for solar panel owners under these two market mechanisms is
@@
Limited supply
$cG\leq L$
prt $\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$ cGcG $L\mathbb{1}\{\overline{\pi}^{\mathrm{b}}_{i}\geq\pi\}$
@@
When supply is abundant, i.e., the total realized solar generation exceeds demand, the resulting prices and allocations are identical under both types of real-time markets. This result follows from the fact that solar supply fully covers demand, rendering utility support unnecessary. As a consequence, the zero marginal cost of solar generation results in a zero equilibrium price. When the solar supply is insufficient to cover the load, the market clearing outcomes of the two market types become different. Since the utility backstop is required in this case, the market price for the single-product real-time market will be $\pi_{\mathrm{u}}$. When solar is traded as a separate product, however, we see that the scarce solar production $cG$ is allocated among the buyers with higher solar premiums first. The solar market price turns out to have a premium over the utility-supplied energy, driven by the competition among the buyers.
@@
and $\pi$ is the solution to $c=\widehat{d}^{\star}(\pi).$ Moreover, the total revenue for solar panel owners under the contract-based market is
@@ V Equilibrium Panel Capacities
Building on the short-term market outcomes derived in the previous section, we next study the investment model to characterize the resulting equilibrium panel investment capacities.
@@
While Lemma 4 enables us to use numerical simulations to compare the equilibrium capacities for each market mechanism, deeper insight comes from a theoretical comparison of the resulting aggregate capacities. To enable a more meaningful comparison, we first define a benchmark capacity as follows.
@@ V-A Benchmark: Social Welfare Optimal Panel Investment
$\max_{c\in\mathbb{R}_{+}}\enspace$ $\widetilde{T}\,\mathbb{E}\left[v^{\star}(c,G)L-\pi_{\mathrm{u}}(L-cG)_{+}\right]-\pi_{0}c.$ (18)
The following result provides an analytical characterization of this capacity.
@@ V-B Analytical Comparison of Equilibrium Capacities
    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}.$ (20)
@@
    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}},$ (22)
@@
    $c^{\mathrm{ne}}_{\mathrm{cb}}-c^{\mathrm{ne}}_{\mathrm{prt}}\geq{\color[rgb]{0,0,0}\beta\epsilon}-O\left(\epsilon^{2}\right),$ (23)
@@
The first interesting result from (20) in Theorem 1 is that the single-product real-time market, in general, leads to an under-investment in solar panels relative to social welfare optimal panel capacity. In fact, with $\epsilon>0$, it is easy to identify a distribution $F_{\widetilde{V}}$ such that $c^{\mathrm{ne}}_{\mathrm{srt}}<c_{\mathrm{opt}}$, and therefore the single-product real-time energy market leads to under-investment. While one could argue that the single-product real-time market is easier to implement, particularly given the limitations of current metering infrastructure, it results in a loss of social welfare when viewed through the lens of long-term investment equilibrium.
The second observation form (20) is that the product-differentiated real-time market is guaranteed to achieve the socially optimal panel investment. This stems from the fact that the product-differentiated market indeed allocates solar energy in the same way as the social welfare optimal allocation in (17). This leads to the same expected revenue for solar panel owners and therefore the same condition for the resulting aggregate panel capacities.
@@ VI Heterogeneous Operation Periods
So far, we have assumed that the values/distributions of key parameters, such as loads and solar irradiation, are identical across operation periods. In practice, however, these values/distributions may vary over time due to factors such as seasonality, time of day, and day of the week. In this section, we extend our analysis to incorporate this temporal heterogeneity.
In our extended model, each period $t$ is characterized by a distinct set of parameters: the total consumer load $L_{t}$, the solar generation per unit panel $G_{t}$ with associated probability density function $f_{G_{t}}$, and the utility backstop price $\pi_{\mathrm{u}_{t}}$. Under this setup, real-time markets are cleared independently for each $t\in\mathcal{T}$, and the expected revenues of solar panel investors are determined accordingly. In contrast, contract-based markets involve long-term agreements, where consumers rent panels from suppliers once for the entire $\mathcal{T}$.
Unlike the homogeneous case, where the expected revenue is identical across periods, here the revenues vary across operation periods. Hence, we compute them individually and then take their average. As in the homogeneous case, we then scale this average by $\widetilde{T}$ to obtain the expected total revenue $\Pi_{m}^{\mathrm{s}}(c)$.
@@ VI-1 CE of solar markets
$\max_{d_{i}\geq 0}\enspace\sum_{t\in\mathcal{T}}\left[v_{i}\mathbb{E}\min\{d_{i}G_{t},L_{t}\}-\pi_{\mathrm{u}_{t}}\mathbb{E}(L_{t}-d_{i}G_{t})_{+}\right]-\pi d_{i}.$
Similar to (9), for each individual $t\in\mathcal{T}$, we can define $\mu_{G_{t,\mathrm{tr}}}(d_{i}):=\mathbb{E}\left[G_{t}\mathbb{1}\bigl{\{}d_{i}G_{t}\leq L_{t}\bigl{\}}\right]$. Using this we can analytically characterize $d_{i}^{\star}(\pi)$ as a solution to $\sum_{t\in\mathcal{T}}\left(\pi_{\mathrm{u}_{t}}+v_{i}\right)\mu_{G_{t,\mathrm{tr}}}\left(d_{i}\right)=\pi.$ The left-hand side of this equation represents the consumer’s expected total valuation for one unit of rented panel capacity over the entire planning period. We treat this quantity as a monotonically decreasing function, denoted by $w_{i}(d_{i})$, and define an extension of its inverse as
@@ VI-3 Social welfare optimal investment
$\max_{c\in\mathbb{R}_{+}}\enspace$ $\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\!\left[v_{t}^{\star}(c,G_{t})L_{t}\!-\!\pi_{\mathrm{u}_{t}}(L_{t}-cG_{t})_{+}\right]\!-\!\pi_{0}c.$ (28)
@@ VII-A Case Study and Problem Description
We conduct numerical experiments in this section to assess our theoretical results. Given California’s leading role in solar adoption, we focus our analysis on this state and consider a large group of potential investors in late 2023 deciding whether to install solar panels across various locations within it. To simplify the analysis, we assume a single pooled distribution-level market for the entire state and consider the three representative market frameworks discussed earlier.
We consider a 25-year planning horizon, with hourly operation periods. While our analytical results are applicable to any number of heterogeneous operation periods, we simplify the case study by grouping operation periods into two representative categories. The first group corresponds to daytime hours, between sunrise and sunset, when solar irradiation is available. The second group includes the remaining hours, during which solar irradiation is minimal or absent. This setup is equivalent to considering $\mathcal{T}=\{1,2\}$, where $t=1$ corresponds to the first group and $t=2$ to the second, with appropriate weights applied to each group in the expected revenue calculations.
Solar irradiation data from 2000 to 2022 for San Francisco, California, are used to model the distributions $f_{G_{t}}(g)$ for all $t\in\mathcal{T}$ [25]. To account for panels’ efficiency, all irradiation values are scaled by a factor of 0.2 to obtain the effective solar irradiation [26]. A kernel density estimator is then used to fit a probability density function to the gathered data, modeling the daytime irradiation density $f_{G_{1}}(g)$. The resulting density function is evaluated against the empirical distribution obtained from the histogram of the data in Fig. 3. Hours with zero or negligible effective irradiation (below $0.1~\mathrm{W}/\mathrm{m}^{2}$) are grouped into the second operation period, represented by a random variable $G_{2}$ that equals zero with probability 1.
We use survey data from [27] on consumers’ willingness to pay a premium for renewable electricity and convert these values into the per-kWh solar premium $v_{i}$ used in our model. As the data is relative to the average electricity bill for the users, we assume an average monthly household energy consumption of $600~\mathrm{kWh}$ and convert the data into solar premium for $1\,\mathrm{kWh}$ of solar energy. We also adjust the data by applying the compound inflation rate in the U.S. from 1999 to 2023. We then fit a truncated exponential density function to the data to model $f_{V}(v_{i})$, which resulted in a mean of 2.86 ¢\$/kWh and an upper bound of 16.57 ¢\$/kWh. Both $\pi_{\mathrm{u}_{1}}$ and $\pi_{\mathrm{u}_{2}}$ are set at 29 ¢\$/kWh, the average retail electricity price in California in 2023 [28]. Moreover, we set $L_{1}=27\,\mathrm{GWh}$ and $L_{2}=29\,\mathrm{GWh}$ [29]. We consider an average capital and installation cost of \$2.7 per watt [30] of solar panels, which includes labor and equipment costs.

### v2

@@ III-B Real-time Pricing Mechanisms
Within the product-differentiated real-time market, the traded product is solar electricity. Each seller $i\in\mathcal{I}_{\mathrm{s}}$ can sell any quantity of solar generation within the set $\mathcal{Q}_{i}^{\mathrm{s}}=[0,cG]$, which depends on the realized value of $G$. The seller payoff function for the representative operation period is
$\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)=\pi q^{\mathrm{s}}_{i},$ (4)
where market equilibrium price $\pi$ and cleared quantity $q^{\mathrm{s}}_{i}$ will be defined momentarily. Given the inelastic load $L$, a buyer $i\in\mathcal{I}_{\mathrm{b}}$ can purchase any quantity of solar within the range $\mathcal{Q}^{\mathrm{b}}_{i}=[0,L]$. Any remaining demand not met by solar will be fulfilled by the utility supply at the price $\pi_{\mathrm{u}}$. The buyer payoff function for the representative operation period is
$\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=(v_{i}-\pi)q^{\mathrm{b}}_{i}-\pi_{\mathrm{u}}(L-q^{\mathrm{b}}_{i}),$ (5)
where $q_{i}^{\mathrm{b}}$ is the cleared quantity for the buyer. The first term in (5) is the buyer’s surplus in the market, and the second term accounts for the payment for the residual demand served by the utility.
Denote $q^{\mathrm{s}}:=\{q^{\mathrm{s}}_{i}\}_{i\in\mathcal{I}_{\mathrm{s}}}$ and $q^{\mathrm{b}}:=\{q^{\mathrm{b}}_{i}\}_{i\in\mathcal{I}_{\mathrm{b}}}$. We use the following equilibrium notion to characterize the market outcome:
@@
We intentionally omit the price discovery process and focus on the CE, as our interest lies in the outcomes rather than the mechanism used to obtain them. Such outcomes can arise, for instance, from buyers and sellers submitting bids to a market operator who clears the market. In our non-atomic game model, where all agents are infinitesimally small and act as price takers, we show that an NE of such a bidding process corresponds to a CE that we consider here (see Appendix A).
With the equilibrium price $\pi$ and the cleared quantities $q^{\mathrm{s}}_{i}$ for sellers in the product-differentiated real-time market, the total expected revenue for sellers with an aggregate panel capacity $c$ over the panel lifespan can be calculated as
@@ III-C Contract-based Market
Remark 3 (Time-scale of contract-based market)
In this section, we assume contracts cover only a single operation period. Although contracts are often long-term in practice, assuming operation periods are identically distributed implies that market outcomes repeat across periods. Thus, contracting for one representative period yields the same outcomes as contracting for multi-period blocks. Importantly, contracts are still set ex ante. Section VI relaxes this assumption and considers heterogeneous periods and multi-period agreements.
@@
where $(z)_{+}:=\max(z,0)$, and $q_{i}^{\mathrm{s}}$ and $q_{i}^{\mathrm{b}}$ are the cleared quantities (i.e., panel capacities) for the seller and buyer, respectively. In (7), the first two terms capture the buyer’s expected surplus in the market, while the last term is the expected payment to the utility for any residual demand.
@@
The proposed market model serves as a general framework capable of representing a variety of solar market mechanisms, provided that the traded product and the payoff structures for buyers and sellers are appropriately defined such that a CE exists. For example, both real-time pricing and contract-based arrangements can be formulated within this framework, with key distinctions captured through the specification of the traded product, the feasible sets of buyers and sellers, and the corresponding payoff functions. We abstract away additional implementation-specific features, as our primary objective is to investigate how different market structures shape long-term investment outcomes. While these mechanisms may differ in operational frequency, often leading to variations in price volatility as discussed in prior work [9, 10], such factors do not directly affect the investment equilibrium when only expected returns matter.
@@
IV Competitive Equilibria of Short-Term Markets
In this section, we derive the expected revenue for panel owners as a function of panel capacity $c$. Proofs are provided in the appendices.
@@ IV Competitive Equilibria of Short-Term Markets
A CE for product-differentiated real-time market (prt) and single-product real-time market (srt), are listed in Table I, where $\overline{F}_{V}(\cdot):=1-F_{V}(\cdot)$ is the complementary cumulative distribution function of $v_{i}$, and $\overline{\pi}^{\mathrm{b}}_{i}:=\pi_{\mathrm{u}}+v_{i}$. Furthermore, the expected total revenue for panel owners under these two market mechanisms is
@@
Limited supply $cG\leq L$ prt $\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$ cGcG $L\mathbb{1}\{\overline{\pi}^{\mathrm{b}}_{i}\geq\pi\}$
@@
When supply is abundant, i.e., the total realized solar generation exceeds demand, the resulting prices and allocations are identical under both types of real-time markets. This result follows from the fact that solar supply fully covers demand, rendering utility support unnecessary. As a consequence, the zero marginal cost of solar generation results in a zero equilibrium price. When the solar supply is insufficient to cover the load, the market clearing outcomes of the two market types become different. Since the utility supply is required in this case, the market price for the single-product real-time market will be $\pi_{\mathrm{u}}$. When solar is traded as a separate product, however, we see that the scarce solar production $cG$ is allocated among the buyers with higher solar premiums first. The solar market price turns out to have a premium over the utility-supplied energy, driven by the competition among the buyers. Accordingly, substituting $\pi$ and $q_{i}^{\mathrm{s}}$ from Table I into (4) yields individual seller payoffs; aggregating via (6) then gives (9). The expression in (10) follows by a similar calculation under the single-product market model.
@@
and $\pi$ is the solution to $c=\widehat{d}^{\star}(\pi).$ Moreover, the total revenue for panel owners under the contract-based market is
@@ V Equilibrium Panel Capacities
Given the short-term market outcomes derived in the previous section, we next apply the investment model to characterize the resulting equilibrium panel investment capacities.
@@
While Lemma 4 enables us to compare the equilibrium capacities across market mechanisms via numerical simulations, deeper insight comes from a theoretical comparison of the resulting aggregate capacities. To enable a more meaningful comparison, we first define a benchmark capacity as follows.
@@ V-A Benchmark: Social Welfare Optimal Panel Investment
$\max_{c\in\mathbb{R}_{+}}\kern 5.0pt$ $\widetilde{T}\,\mathbb{E}\left[v^{\star}(c,G)L-\pi_{\mathrm{u}}(L-cG)_{+}\right]-\pi_{0}c,$ (20)
where the objective function of (20) sums consumers and investors payoffs; internal monetary transfers between them cancel out. The following result provides an analytical characterization of this capacity.
@@ V-B Analytical Comparison of Equilibrium Capacities
    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}.\vskip-2.5pt$ (22)
@@
    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}},\vskip-2.5pt$ (24)
@@
    $c^{\mathrm{ne}}_{\mathrm{cb}}-c^{\mathrm{ne}}_{\mathrm{prt}}\geq{\color[rgb]{0,0,0}\beta\epsilon}-O\left(\epsilon^{2}\right),\vskip-2.5pt$ (25)
@@
The first interesting result from (22) in Theorem 1 is that the single-product real-time market, in general, leads to an under-investment in panels relative to social welfare optimal panel capacity. In fact, with $\epsilon>0$, it is easy to identify a distribution $F_{\widetilde{V}}$ such that $c^{\mathrm{ne}}_{\mathrm{srt}}<c_{\mathrm{opt}}$. While one could argue that the single-product real-time market is easier to implement, particularly given the limitations of current metering infrastructure, it results in a loss of social welfare when viewed through the lens of long-term investment equilibrium.
The second observation [^6] In this case, an analytical comparison between $c_{\mathrm{prt}}^{\mathrm{ne}}$ and $c_{\mathrm{cb}}^{\mathrm{ne}}$ is not available; nevertheless, our numerical results indicate that $c_{\mathrm{cb}}^{\mathrm{ne}}>c_{\mathrm{prt}}^{\mathrm{ne}}$ in most realistic setups. from (22) is that the product-differentiated real-time market is guaranteed to achieve the socially optimal panel investment. This stems from the fact that the product-differentiated market indeed allocates solar energy in the same way as the social welfare optimal allocation in (19). This leads to the same expected revenue for panel owners and therefore the same condition for the resulting aggregate panel capacities. Nonetheless, implementing such a market in practice is nontrivial and would require granular metering, verification, and settlement to track solar deliveries.
@@ VI Heterogeneous Operation Periods
So far, we have assumed that the values/distributions of key parameters, such as loads and solar irradiation, are identical across operation periods. In practice, however, these values/distributions may vary over time (e.g., by season, time of day, and day of week). In this section, we extend our analysis to incorporate this temporal heterogeneity.
In our extended model, each period $t$ is characterized by a distinct set of parameters: the total consumer load $L_{t}$, the solar generation per unit panel $G_{t}$ with associated probability density function $f_{G_{t}}$, and the utility supply price $\pi_{\mathrm{u}_{t}}$. In this model, real-time markets are cleared independently for each $t\in\mathcal{T}$, and the expected revenues of panel investors are determined accordingly. In contrast, contract-based markets involve long-term agreements, where consumers rent panels from suppliers once for the entire $\mathcal{T}$. Unlike the homogeneous case, where expected revenue is identical across periods, revenues here vary by operation period. We therefore compute period-by-period revenues, average them, and then scale by $\widetilde{T}$ to obtain the expected total revenue $\Pi_{m}^{\mathrm{s}}(c)$.
@@ VI-1 CE of solar markets
$\max_{d_{i}\geq 0}\!\!\kern 5.0pt\sum_{t\in\mathcal{T}}\left[v_{i}\mathbb{E}\min\{d_{i}G_{t},L_{t}\}\!-\!\pi_{\mathrm{u}_{t}}\mathbb{E}(L_{t}-d_{i}G_{t})_{+}\right]\!-\!\pi d_{i}.$ (29)
Similar to (11), for each individual $t\in\mathcal{T}$, we can define $\mu_{G_{t,\mathrm{tr}}}(d_{i}):=\mathbb{E}\left[G_{t}\mathbb{1}\bigl\{d_{i}G_{t}\leq L_{t}\bigl\}\right]$. Using this we can analytically characterize $d_{i}^{\star}(\pi)$ as a solution to $\sum_{t\in\mathcal{T}}\left(\pi_{\mathrm{u}_{t}}+v_{i}\right)\mu_{G_{t,\mathrm{tr}}}\left(d_{i}\right)=\pi.$ The left-hand side of this equation represents the consumer’s expected total valuation for one unit of rented panel capacity over the entire planning period. We treat this quantity as a non-increasing function, denoted by $w_{i}(d_{i})$, and define an extension of its inverse as
@@ VI-3 Social welfare optimal investment
$\max_{c\in\mathbb{R}_{+}}\kern 5.0pt$ $\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\!\left[v_{t}^{\star}(c,G_{t})L_{t}\!-\!\pi_{\mathrm{u}_{t}}(L_{t}-cG_{t})_{+}\right]\!-\!\pi_{0}c.$ (31)
@@ VII-A Case Study and Problem Description
In this section, we conduct numerical experiments to evaluate our theoretical results. Motivated by California’s leading solar adoption, we model a large population of potential investors in late 2023 deciding whether to install PV across locations statewide. To simplify the analysis, we assume a single statewide pooled distribution-level market and consider the three representative market frameworks introduced earlier.
We consider a 25-year panel lifespan with hourly operation periods. Although our analysis allows for arbitrarily many heterogeneous periods, we group hours into two representative categories: daytime (sunrise-sunset) with solar irradiation, and nighttime with minimal or no irradiation. Equivalently, we set $\mathcal{T}=\{1,2\}$, where $t=1$ denotes daytime and $t=2$ nighttime, and weight each group in the expected-revenue calculations by its relative frequency (i.e., the number of hours it represents).
We use solar irradiation data from 2000 to 2022 for San Francisco, California, to model $f_{G_{t}}(g)$ for all $t\in\mathcal{T}$ [24]. To account for panel efficiency, we scale irradiation by a factor of $0.2$ to obtain the effective solar irradiation [25]. A kernel density estimator is then used to fit a probability density function to the gathered data, modeling the daytime irradiation density $f_{G_{1}}(g)$. The resulting density function is evaluated against the empirical distribution obtained from the histogram of the data in Fig. 3. Hours with zero or negligible effective irradiation (below $0.1~\mathrm{W}/\mathrm{m}^{2}$) are grouped into the second operation period, represented by a random variable $G_{2}$ that equals zero with probability 1.
We use survey data from [23] on consumers’ willingness to pay a premium for renewable electricity and convert these values into the per-kWh solar premium $v_{i}$ used in our model. As the data is relative to the average electricity bill for the users, we assume an average monthly household energy consumption of $600~\mathrm{kWh}$ and convert the data into solar premium for $1\,\mathrm{kWh}$ of solar energy. We also adjust the data by applying the compound inflation rate in the U.S. from 1999 to 2023. We then fit a truncated exponential density function to the data to model $f_{V}(v_{i})$, which resulted in a mean of 2.86 ¢\$/kWh and an upper bound of 16.57 ¢\$/kWh. Both $\pi_{\mathrm{u}_{1}}$ and $\pi_{\mathrm{u}_{2}}$ are set at 29 ¢\$/kWh, the average retail electricity price in California in 2023 [26]. Moreover, we set $L_{1}=27\,\mathrm{GWh}$ and $L_{2}=29\,\mathrm{GWh}$ [27]. We consider an average capital and installation cost of \$2.7 per watt [28] for solar panels, which includes labor and equipment costs.